OpenStudy (anonymous):

The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

pooja195 (pooja195):

Surface Area = 2Lw + 2wh + 2Lh = 150cm² where: L = length w = width h = height So, if you new sides are trippled: Surface Area = 2(3L)(3w) + 2(3w)(3h) + 2(3L)(3h) = 18Lw + 18wh + 18Lh Set up the ratio of new surface area to old: (18Lw + 18wh + 18Lh)/(2Lw + 2wh + 2Lh) = Area/150 and solve for Area.

pooja195 (pooja195):

then divide

pooja195 (pooja195):

no 150 didu read above

OpenStudy (anonymous):

2.77777

pooja195 (pooja195):

round it and get 2.7 dou have answer choices

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

its an explain question

pooja195 (pooja195):

ok then lets go with 2.7

OpenStudy (anonymous):

alright 2.7cm

OpenStudy (anonymous):

right 2.7cm?

pooja195 (pooja195):

yep

OpenStudy (anonymous):

how about this problem The volume of a rectangular prism is 183 in³. What is the volume of a pyramid which has the same height and a congruent base?

pooja195 (pooja195):

Prism Volume = Lwh Pyramid Volume = 1/3(Lwh)

OpenStudy (anonymous):

1/3 (183)?

pooja195 (pooja195):

yes thats rite

OpenStudy (anonymous):

i got 61in^3

OpenStudy (anonymous):

Find the surface area of a small juice can whose diameter is 4 inches and whose height is 6.25 inches. Use 3.14 for π. Hint: The can is a cylinder.

pooja195 (pooja195):

Surface Area of a Cylinder = 2πr² + 2πrh where r = radius: d/2 h = height

OpenStudy (anonymous):

ok os what do i do now

OpenStudy (anonymous):

2(3.14) * 2(3.14)?

pooja195 (pooja195):

yes an when u get the answer message me srry im in high demand i have 10 more people to help

OpenStudy (anonymous):

i got 39.4384

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

correct

pooja195 (pooja195):

yes.

OpenStudy (anonymous):

The volume of a right circle cone is 800 mm³ and its radius is 5 mm. Find the height of the cone. Use 3.14 for π.

pooja195 (pooja195):

base area = radius² x 3.14 base area = 8² x 3.14 base area = 64 x 3.14 base area = 200.96 volume = base area x height : 3 height = 3 x volume : base area height = 3 x 500 : 200.96 height = 7.46

pooja195 (pooja195):

mm

pooja195 (pooja195):

area = 4 π r² area = 4 π 0.25² area = 4 π 0.0625 area = 0.25 π

pooja195 (pooja195):

yep good job